# Pangu DeepDiver: Adaptive Search Intensity Scaling via Open-Web Reinforcement Learning

### v1

@@ 5.4 Generalization to Open-ended Problems
Setup We evaluate DeepDiver on ProxyQA (Tan et al., 2024) against R1-Distilled baselines. Since DeepDiver generates Chinese responses, we translate all ProxyQA meta-questions and sub-questions for evaluation. Testing prompt and evaluator configuration follow the original study. We analyze generalization benefits gained through RL training compared to distillation.
ProxyQA
Acc. # rounds # searches response length
R1-Distill 23.25% 1.76 7.39 590.31
DeepDiver (Ours) 32.72% 2.27 10.54 1971.58
Results RL training significantly enhances the generalization capability of LLMs, enabling transition from closed-ended to open-ended problems and demonstrating strong adaptability to long-form writing tasks. As shown in Table 4, our RL-guided DeepDiver achieves $32.72\%$, outperforming the R1-distilled model by $9.47$ percentage points. This suggests RL training enables more effective information seeking and validation in open-web environments, resulting in more comprehensive responses. Additionally, DeepDiver’s response length and search queries are substantially higher than the distilled model’s, providing evidence that search intensity scaling encourages active information acquisition for more comprehensive answers. We show two examples without cherry picking in Figure 7, Figure 8, Figure 9 and Figure 10 to show the obvious differences between our DeepDiver and distilled baselines.
5.5 Search Intensity vs. Difficulty
@@ Results
5.6 Statistics of Information-Seeking Behaviors
To further analyze the differences between our WebPuzzle dataset and wiki-based datasets, as well as to conduct an in-depth investigation into the information seeking behavior of LLMs, we computed the proportions of different information-seeking behaviors (defined in Section 2.2) across various models on multiple datasets. The detailed results are presented in Table 6.
@@ 5.6 Statistics of Information-Seeking Behaviors
Following the setting of  Gandhi et al. (2025), we developed an automatic pipeline based on GPT-4o (OpenAI, 2024b) to identify and count the occurrences of different behaviors in the reasoning chains of model outputs. We primarily focused on three behaviors: Reflection & Correction, Conflict Resolution, and Verification & Denoising, while omitting the statistics for Evidence Gathering & Supplements due to its overly fundamental nature. Detailed prompt specifications can be found in Figure 15. The baseline models, evaluation benchmarks, and analysis setup align with those in Section 5.2.
@@ Results
Our proposed WebPuzzle benchmark is more challenging than existing wiki-based datasets, requiring more sophisticated information-seeking behaviors. Furthermore, DeepDiver trained with a real-world web environment and problems exhibits more diverse and complex information-seeking behavior compared to other methods. As shown in Table 6, when evaluated on WebPuzzle, all baseline models demonstrate more complex reasoning patterns than on wiki-based datasets. This highlights WebPuzzle’s ability to present more intricate and demanding tasks than prior benchmarks. Moreover, from the perspective of model behavior, DeepDiver consistently generates significantly richer information-seeking trajectories than other baselines. These results underscore the benefits of training in real-world web environments, both in terms of the complexity of the tasks and the realism of interaction with real-world search engines, outperforming models trained on wiki-based corpora.
5.7 Human vs. DeepDiver
@@
Five human experts, who were not involved in the annotation of the evaluation sets, participated in the human evaluation. Each expert was tasked with answering 5 problems, creating a subset of the evaluation set. During testing, the number of search queries and web pages browsed was recorded. Consistent with the experimental setup, we report the average accuracy of the human evaluators and compare it against DeepSeek-R1 and our proposed DeepDiver to highlight the challenges posed by WebPuzzle.
@@
WebPuzzle presents significant challenges, even for human experts, who face difficulties with numerous searches and reasoning steps. As shown in Table 7, human evaluators achieved an accuracy rate of 44.0%, requiring an average of 6.16 search queries and browsing 9.28 web pages to solve the problems. In comparison, DeepSeek-R1 made 7.37 search queries with 1.94 rounds of searching. Despite DeepSeek-R1 performing the most search rounds among all baselines and humans, it achieved a relatively lower accuracy of 38.7%. In contrast, our proposed DeepDiver conducted 10.69 search queries across 2.68 rounds, significantly surpassing the other methods in terms of search effort. This additional effort allowed DeepDiver to more thoroughly collect and verify evidence, leading to a 40.0% accuracy, which is closer to human performance. These results suggest that DeepDiver follows a more comprehensive information-seeking process, better aligning with the approach taken by human evaluators.
5.8 Case Study
@@
DeepDiver demonstrates exceptional information-seeking ability while incorporating correction on reasoning history and retrieved documents, providing a more robust and adaptable solution for overcoming the limitations of flawed internal knowledge. Specifically, as shown in Table D.1, DeepSeek-R1 leverages its rich internal knowledge to quickly narrow the exploration scope, consistently demonstrating the ability to list the correct answer among candidate options in the first round, showcasing remarkable knowledge retention. Taking advantage of this, R1 can focus more on verifying whether candidate answers satisfy all constraints when designing search queries, allowing it to find the correct answer in fewer rounds. However, the R1-distilled Qwen-7B attempts to mimic DeepSeek-R1’s behavior but lacks error correction when internal knowledge is flawed. Specifically, the R1-distilled Qwen suggests “Nico Hülkenberg” in the first round of searching and reasoning, but fails to resolve conflicting conditions in subsequent rounds due to limited internal knowledge, ultimately producing a faulty answer.
@@
6 Related Work
Prompting-based strategies—including in-context learning (Brown et al., 2020) and retrieval-augmented chain-of-thought pipelines (Jiang et al., 2023; Trivedi et al., 2023; Shao et al., 2023)—enable zero- or few-shot question answering, yet their fixed templates rarely adapt retrieval depth to unforeseen information gaps. Supervised fine-tuning (SFT) improves the synergy between retrieval and generation (Asai et al., 2023; Yu et al., 2024) but can overfit corpus-specific inference patterns, hindering transfer to noisy settings. Reinforcement-learning (RL) methods let LLMs decide when and what to search, achieving state-of-the-art results on curated benchmarks such as HotpotQA (Jin et al., 2025; Song et al., 2025; Chen et al., 2025a; Zheng et al., 2025; Yang et al., 2018; Sun et al., 2025), yet they remain evaluated mostly in “clean” Wikipedia-style environments. Beyond these directions, tool-augmented agents that interleave reasoning with web search (Nakano et al., 2022; Yao et al., 2023; Yang et al., 2023) similarly demonstrate promise but still rely on limited test beds, underscoring the need for benchmarks that reflect real-world, noisy information-seeking scenarios. Additional introduction of the related works is shown in Appendix A.
7 Discussion: Limitations and Extensions
@@
We demonstrate that DeepDiver, trained on closed-ended open-web problems, exhibits strong generalization capabilities when applied to open-ended tasks, such as long-form writing. However, developing a more effective RL-driven framework that directly enhances the model’s ability to solve these problems remains an open challenge. Since the RL framework relies on a stable reward signal, the lack of reliable metrics to assess open-ended content complicates the task of defining such signals. As a result, designing a framework that can handle both open-ended and closed-ended problems remains a key area for future exploration. One potential approach could be following the approaches of ProxyQA, involving crafting sub-questions in an online way during RL training, which may offer a promising direction for enhancing the model’s performance on open-ended tasks.
@@
Search engines are commonly viewed as tools that LLMs can utilize to enhance their capabilities. In our study, we focused solely on investigating the reasoning and searching behavior of LLMs with the aid of search engines. This limits the scope of our work, as there are various other tools, such as those compatible with the Model Context Protocol (MCP) Anthropic (2024), which could also contribute to improving reasoning and searching processes in knowledge-intensive tasks. Future research could expand on this by considering the integration of additional tools to further enhance the performance of LLMs in such tasks.
@@
Prior work has shown that reasoning LLMs trained in RL environments often suffer from “overthinking” (Chen et al., 2025b), generating excessively long reasoning sequences—sometimes spanning thousands of tokens—even for simple questions. We observe a similar phenomenon in our DeepDiver model, which invokes significantly more search calls compared to other baselines, even when evaluated on simple tasks. This over-searching behavior highlights an inefficiency that future work should aim to mitigate. A promising direction would be to develop methods that reduce both the search and reasoning overhead in LLM-based systems.
8 Conclusion
Appendix
Appendix A Additional Related Works
Datasets and Benchmarks
Existing datasets for evaluating LLMs’ information-seeking abilities, such as HotpotQA (Yang et al., 2018) and Bamboogle (Press et al., 2023), focus on multi-hop QA. HotpotQA offers 113k Wikipedia-based questions with sentence-level rationales, while Bamboogle provides 125 hand-crafted two-hop questions whose answers cannot be surfaced directly by search engines. However, their reliance on Wikipedia limits the adaptation to real-world problems—modern LLMs often memorize the wiki-based contents, and wiki-based search environments are too clean to reflect real-world web conditions. Concurrent to our work, recent benchmarks like BrowseComp (Wei et al., 2025a) and BrowseComp-zh (Zhou et al., 2025) push information seeking evaluation further by requiring multiple searches and deep link navigation in an open-web environment. Yet, they only offer evaluation data. In contrast, our proposed WebPuzzle dataset includes 24k training examples spanning riddle-style and cross-page QA, better enabling the information seeking ability of the LLMs.
Prompting-based strategies.
Few-shot in-context learning (Brown et al., 2020) allows frozen LLMs to imitate reasoning patterns from exemplars, while RAG with CoT prompts (Jiang et al., 2023; Trivedi et al., 2023; Shao et al., 2023) interleave search queries with intermediate thoughts to inject fresh evidence. These methods require no training, but their step counts and query formats are anchored to the prompt, so the model cannot escalate effort when initial evidence is missing. As a result, they often stop searching too early or hallucinate unsupported facts on hard, open-web questions.
Supervised fine-tuning (SFT).
Retrieval-augmented generation systems fine-tuned on gold passages—e.g., Self-RAG and Auto-RAG (Asai et al., 2023; Yu et al., 2024)—learn to quote and merge the external text into answers, reducing hallucinations on Wikipedia benchmarks. Nevertheless, SFT tends to overfit the inference paradigm tied to the training corpus; performance drops when pages are noisy, multilingual, or partially missing, which is common in the open-web internet environment.
Reinforcement learning (RL).
RL integrated iterative RAG pipeline offers a principled way for LLM agents to decide when and what to search. Recent work trains search-capable agents with reward shaping and curricula, achieving state-of-the-art accuracy on HotpotQA and similar wiki-based corpus (Jin et al., 2025; Song et al., 2025; Chen et al., 2025a; Zheng et al., 2025). Yet these studies use closed Wikipedia environments, where every answer is guaranteed to exist and pages are clean.
Tool-augmented agents.
ReAct-style frameworks combine chain-of-thought with executable actions, letting models call a browser or other tools between reasoning steps (Nakano et al., 2022; Yao et al., 2023; Yang et al., 2023; Wei et al., 2025b). They excel at citing fresh evidence and correcting themselves mid-trajectory, but they still struggle with irrelevant pages and hard to deal with the real-world open-web environment.
Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a reinforcement learning algorithm designed to enhance the efficiency of Proximal Policy Optimization (PPO) (Schulman et al., 2017) by eliminating the need for a critic network. Specifically, GRPO samples multiple outputs from a previous policy for a given prompt and computes their average reward to serve as a dynamic baseline. The advantage of each output is defined relative to this baseline, resulting in positive advantages for outputs exceeding the baseline and negative advantages otherwise. Formally, we define the GRPO objective as follows:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}\left[\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(o_{i})}{\pi_{\theta_{\text{old}}}(o_{i})}A_{i},\;\text{clip}\left(\frac{\pi_{\theta}(o_{i})}{\pi_{\theta_{\text{old}}}(o_{i})},1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta\,\mathbb{D}_{\text{KL}}\left(\pi_{\theta}\,\|\,\pi_{\text{ref}}\right)\right],$ (10)
where the relative advantage $A_{i}$ is computed as:
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\dots,r_{G}\})}.$ (11)
Appendix B Data Curation
B.1 Principles to Annotate the Evaluation Set
@@
B.2 Quality Assurance of WebPuzzle
We prioritize recently updated pages not fully covered by LLMs’ knowledge cutoffs, emphasizing open-web searches for problem-solving. We exclude offensive, politically sensitive, ethically concerning, or NSFW content using an LLM-based filter. Ambiguous, controversial, multiple-choice, and boolean questions are removed to prevent answer hacking. We also eliminated unsolvable problems to ensure low error rates, ultimately collecting 24k WebPuzzle training samples.
Appendix C Detailed Experimental Setup
C.1 Reward Definition

### v2

@@ 5.5 Generalization to Open-ended Problems
Setup We evaluate DeepDiver on ProxyQA [tan2024proxyqaalternativeframeworkevaluating] against R1-Distilled baselines. Since DeepDiver generates Chinese responses, we translate all ProxyQA meta-questions and sub-questions for evaluation. Testing prompt and evaluator configuration follow the original study. We analyze generalization benefits gained through RL training compared to distillation.
Results RL training significantly enhances the generalization capability of LLMs, enabling transition from closed-ended to open-ended problems and demonstrating strong adaptability to long-form writing tasks. As shown in Table 6, our RL-guided DeepDiver achieves $32.72\%$, outperforming the R1-distilled model by $9.47$ percentage points. This suggests RL training enables more effective information seeking and validation in open-web environments, resulting in more comprehensive responses. Additionally, DeepDiver’s response length and search queries are substantially higher than the distilled model’s, providing evidence that search intensity scaling encourages active information acquisition for more comprehensive answers.
6 Related Work
Prompting-based strategies—including in-context learning [brown2020languagemodelsfewshotlearners] and retrieval-augmented chain-of-thought pipelines [jiang2023activeretrievalaugmentedgeneration, trivedi2023interleavingretrievalchainofthoughtreasoning, shao2023enhancingretrievalaugmentedlargelanguage]—enable zero- or few-shot question answering, yet their fixed templates rarely adapt retrieval depth to unforeseen information gaps. Supervised fine-tuning (SFT) improves the synergy between retrieval and generation [asai2023selfraglearningretrievegenerate, yu2024autoragautonomousretrievalaugmentedgeneration] but can overfit corpus-specific inference patterns, hindering transfer to noisy settings. Reinforcement-learning (RL) methods let LLMs decide when and what to search, achieving state-of-the-art results on curated benchmarks such as HotpotQA [jin2025search, song2025r1searcherincentivizingsearchcapability, chen2025research, zheng2025deepresearcher, yang2018hotpotqadatasetdiverseexplainable], yet they remain evaluated mostly in “clean” Wikipedia-style environments. Beyond these directions, tool-augmented agents that interleave reasoning with web search [nakano2022webgptbrowserassistedquestionansweringhuman, yao2023reactsynergizingreasoningacting, yang2023gpt4toolsteachinglargelanguage] similarly demonstrate promise but still rely on limited test beds, underscoring the need for benchmarks that reflect real-world, noisy information-seeking scenarios. Additional introduction of the related works are shown in Appendix B.
7 Conclusion
Appendix
Appendix A Further Analysis
A.1 Search Intensity vs. Difficulty
@@ Results
A.2 Statistics of Information-Seeking Behaviors
To further analyze the differences between our WebPuzzle dataset and wiki-based datasets, as well as to conduct an in-depth investigation into the behavior of our DeepDiver model, we computed the proportions of different information-seeking behaviors (defined in section 2.2) across various models on multiple datasets. The detailed results are presented in Table 4.
@@ A.2 Statistics of Information-Seeking Behaviors
We developed an automatic pipeline based on GPT-4o [openai_gpt4o] to identify and count the occurrences of different behaviors in the reasoning chains of model outputs. We primarily focused on three behaviors: Reflection & Correction, Conflict Resolution, and Verification & Denoising, while omitting the statistics for Evidence Gathering & Supplements due to its overly fundamental nature. Detailed prompt specifications can be found in Figure 16. The baseline models, evaluation benchmarks, and analysis setup align with those in Experiment 5.2.
@@ Results
The WebPuzzle proves to be more challenging compared to wiki-based datasets, requiring more complex information-seeking behaviors. Additionally, DeepDiver’s reasoning chains exhibit richer and more diverse patterns. As shown in tabel 4, The behavioral statistics of WebPuzzle-en surpass those of wiki-based datasets across nearly all models, particularly in Reflection & Correction and Conflict Resolution behaviors. Notably, for the HotpotQA dataset - the primary training data for both baseline models - we observe a substantial reduction in the proportion of complex behaviors required compared to WebPuzzle when evaluating the same model, further highlighting the importance of more open-web training data. Additionally, DeepDiver, trained on WebPuzzle with real search engine integration, demonstrates richer and more sophisticated information-seeking behaviors compared to various baseline models.
A.3 Human vs. DeepDiver
@@
Five human experts, who were not involved in the annotation of the evaluation sets, participated in the human evaluation. Each expert was tasked with answering 5 problems, creating a subset of the evaluation set. During testing, the number of search queries and web pages browsed were recorded. Consistent with the experimental setup, we report the average accuracy of the human evaluators and compare it against DeepSeek-R1 and our proposed DeepDiver to highlight the challenges posed by WebPuzzle.
@@
WebPuzzle presents significant challenges, even for human experts, who face difficulties with numerous searches and reasoning steps. As shown in Table 5, human evaluators achieved an accuracy rate of 44.0%, requiring an average of 6.16 search queries and browsing 9.28 web pages to solve the problems. In comparison, DeepSeek-R1 made 7.37 search queries with 1.94 rounds of searching. Despite DeepSeek-R1 performing the most search rounds among all baselines and humans, it achieved a relative lower accuracy of 38.7%. In contrast, our proposed DeepDiver conducted 10.69 search queries across 2.68 rounds, significantly surpassing the other methods in terms of search effort. This additional effort allowed DeepDiver to more thoroughly collect and verify evidence, leading to a 40.0% accuracy rate, which is closer to human performance. These results suggest that DeepDiver follows a more comprehensive information-seeking process, better aligning with the approach taken by human evaluators.
A.4 Case Study
@@
DeepDiver demonstrates exceptional information-seeking ability while incorporating correction on reasoning history and retrieved documents, providing a more robust and adaptable solution for overcoming the limitations of flawed internal knowledge. Specifically, as shown in Table 10, DeepSeek-R1 leverages its rich internal knowledge to quickly narrow the exploration scope, consistently demonstrating the ability to list the correct answer among candidate options in the first round, showcasing remarkable knowledge retention. Taking advantage of this, R1 can focus more on verifying whether candidate answers satisfy all constraints when designing search queries, allowing it to find the correct answer in fewer rounds. However, the R1-distilled Qwen-7B attempts to mimic DeepSeek-R1’s behavior but lacks error correction when internal knowledge is flawed. Specifically, the R1-distilled Qwen suggests “Nico Hülkenberg” in the first round of searching and reasoning, but fails to resolve conflicting conditions in subsequent rounds due to limited internal knowledge, ultimately producing a faulty answer.
@@
A.5 Results on ProxyQA
We evaluate DeepDiver on ProxyQA [tan2024proxyqaalternativeframeworkevaluating] against R1-Distilled baselines. Since DeepDiver generates Chinese responses, we translate all ProxyQA meta-questions and sub-questions for evaluation. Testing prompt and evaluator configuration follow the original study. We analyze generalization benefits gained through RL training compared to distillation.
ProxyQA
Acc. # rounds # searches response length
R1-Distill 23.25% 1.76 7.39 590.31
DeepDiver 32.72% 2.27 10.54 1971.58
Appendix B Additional Related Works
Prompting-based strategies.
Few-shot in-context learning [brown2020languagemodelsfewshotlearners] allows frozen LLMs to imitate reasoning patterns from exemplars, while RAG with CoT prompts [jiang2023activeretrievalaugmentedgeneration, trivedi2023interleavingretrievalchainofthoughtreasoning, shao2023enhancingretrievalaugmentedlargelanguage] interleave search queries with intermediate thoughts to inject fresh evidence. These methods require no training, but their step counts and query formats are anchored to the prompt, so the model cannot escalate effort when initial evidence is missing. As a result, they often stop searching too early or hallucinate unsupported facts on hard, open-web questions.
Supervised fine-tuning (SFT).
Retrieval-augmented generation systems fine-tuned on gold passages—e.g., Self-RAG and Auto-RAG [asai2023selfraglearningretrievegenerate, yu2024autoragautonomousretrievalaugmentedgeneration]—learn to quote and merge the external text into answers, reducing hallucinations on Wikipedia benchmarks. Nevertheless, SFT tends to overfit the inference paradigm tied to the training corpus; performance drops when pages are noisy, multilingual, or partially missing, which is common in open-web internet environment.
Reinforcement learning (RL).
RL integrated iterative RAG pipeline offers a principled way for LLM agents to decide when and what to search. Recent work trains search-capable agents with reward shaping and curricula, achieving state-of-the-art accuracy on HotpotQA and similar wiki-based corpus [jin2025search, song2025r1searcherincentivizingsearchcapability, chen2025research, zheng2025deepresearcher]. Yet these studies use closed Wikipedia environments, where every answer is guaranteed to exist and pages are clean.
Tool-augmented agents.
ReAct-style frameworks combine chain-of-thought with executable actions, letting models call a browser or other tools between reasoning steps [nakano2022webgptbrowserassistedquestionansweringhuman, yao2023reactsynergizingreasoningacting, yang2023gpt4toolsteachinglargelanguage]. They excel at citing fresh evidence and correcting themselves mid-trajectory, but they still struggle with irrelevant pages and and hard to dealing with the real-world open-web environment.
Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical] is a reinforcement learning algorithm designed to enhance the efficiency of Proximal Policy Optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms] by eliminating the need for critic network. Specifically, GRPO samples multiple outputs from a previous policy for a given prompt and computes their average reward to serve as a dynamic baseline. The advantage of each output is defined relative to this baseline, resulting in positive advantages for outputs exceeding the baseline and negative advantages otherwise. Formally, we define the GRPO objective as follow:
[…]
where the relative advantage $A_{i}$ is computed as:
$\footnotesize{A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\dots,r_{G}\})}.}$ (2)
Appendix C Discussion: Limitations and Extensions
@@
We demonstrate that DeepDiver, trained on closed-ended open-web problems, exhibits strong generalization capabilities when applied to open-ended tasks, such as long-form writing. However, developing a more effective RL-driven framework that directly enhances the model’s ability to solve these problems remains an open challenge. Since the RL framework relies on a stable reward signal, the lack of reliable metrics to assess open-ended content complicates the task of defining such signals. As a result, designing a framework that can handle both open-ended and closed-ended problems remains a key area for future exploration. One potential approach could following the approaches of ProxyQA, involving crafting sub-questions in an online way during RL training, which may offer a promising direction for enhancing the model’s performance on open-ended tasks.
@@
Search engines are commonly viewed as tools that LLMs can utilize to enhance their capabilities. In our study, we focused solely on investigating the reasoning and searching behavior of LLMs with the aid of search engines. This limits the scope of our work, as there are various other tools, such as those compatible with the Model Context Protocol (MCP) [anthropic2024mcp], which could also contribute to improving reasoning and searching processes in knowledge-intensive tasks. Future research could expand on this by considering the integration of additional tools to further enhance the performance of LLMs in such tasks.
@@
Prior work has shown that reasoning LLMs trained in RL environments often suffer from “overthinking” [chen2025think23overthinkingo1like], generating excessively long reasoning sequences—sometimes spanning thousands of tokens—even for simple questions. We observe a similar phenomenon in our DeepDiver model, which invokes significantly more search calls compared to other baselines, even when evaluated on simple tasks. This over-searching behavior highlights an inefficiency that future work should aim to mitigate. A promising direction would be to develop methods that reduce both the search and reasoning overhead in LLM-based systems.
Appendix D Detailed Dataset Curation
D.1 Examples of the WebPuzzle
In this section, we show two examples of the WebPuzzle sampled from the cross-page question and riddles respectively.
Example of Cross-page Question 2024年10月31日,一名身高125cm的普通成人游客购买深圳欢乐谷万圣节夜场票,需支付多少钱? Translation: On October 31st, 2024, how much does an ordinary adult tourist with a height of 125 cm need to pay for the night ticket of the Halloween event at Happy Valley Shenzhen?   Solution: 149 RMB
Example of Riddle 某中国社交媒体应用因美国政府的禁令而意外获得大量关注,用户自称为“难民”并迁移到该平台。该应用在苹果应用商店一度下载量排名第一,超过ChatGPT。其母公司总部位于上海。这是哪个应用? Translation: A Chinese social media app unexpectedly gained a great deal of attention due to a U.S. government ban. Users refer to themselves as “refugees” and have migrated to this platform. At one point, the app ranked first in downloads on the Apple App Store, surpassing ChatGPT. Its parent company is headquartered in Shanghai. Which app is this?   Solution: 小红书 / Rednote / Xiaohongshu
D.2 Principles to Annotate the Evaluation Set
@@
D.3 Quality Assurance of WebPuzzle
We prioritize recently updated pages not fully covered by LLMs’ knowledge cutoffs, emphasizing open-web searches for problem-solving. We exclude offensive, politically sensitive, ethically concerning, or NSFW content using an LLM-based filter. Ambiguous, controversial, multiple-choice and boolean questions are removed to prevent answer hacking. We also eliminated unsolvable problems to ensure low error rates, ultimately collecting 24k WebPuzzle training samples.
D.4 Diffculty Level Tagging
Formally, let $N_{\text{correct}}$ denote the number of correct answers out of the 4 tests. The difficulty level $D$ of a problem can be determined as:
$\footnotesize{D=\begin{cases}\text{easy}&\text{if }N_{\text{correct}}=4\\ \text{medium}&\text{if }N_{\text{correct}}=2\text{ or }3\\ \text{hard}&\text{if }N_{\text{correct}}=1\\ \text{outlier}&\text{if }N_{\text{correct}}=0\end{cases}}$ (3)
Appendix E Detailed Experimental Setup
E.1 Statistic of WebPuzzle
Data Category Training Data Num Evaluation Data Num
Easy Medium Hard Outliers ALL Easy Medium Hard Outliers ALL
Cross-Page QA 2553 2451 1404 3970 4500 5 54 27 44 130
Open&Wiki Riddle 5566 2956 1409 3375 2500 0 37 33 75 145
Total Set 8119 5407 2813 7345 23684 5 91 60 119 275
E.2 Reward Defination